OpenStudy (anonymous):

Perform the indicated operation. Please show all of your work 2a/3 / 1/2

OpenStudy (dape):

\[ \frac{\frac{2a}{3}}{\frac{1}{2}} = \frac{2*\frac{2a}{3}}{2*\frac{1}{2}} = \frac{\frac{4a}{3}}{1} = \frac{4a}{3} \]